# Pushing Trade-Off Boundaries: Compact yet Effective Remote Sensing Change Detection

### v1

@@ 3.3 Local-Global Fusion Block
After obtaining the difference map $D_{i}$ between the two temporal feature maps, it is essential to further capture the relational information among the changed regions. To this end, we extract relationship features at both local and global scales, motivated by the following observations:
  • On the local scale, when a specific region is surronded by other changed areas, it is highly likely that this region has also undergone change. Modeling such local relationships helps mitigate the presence of hols or fragmentation within detected changed regions.
  • On the global scale, some images contain multiple instances of similar changes, for example, repeated structural changes in buildings. Capturing global contextual relationships enables the model to identify and enhance the consistency among these semantically similar change patterns, thereby improving the extraction accuracy for such repeated changes across the image.
To achieve the aforementioned objectives, we propose the Local-Global Fusion Blick(LGFB). This module consists of two key components: Sliding-Window Self-Attention(SWSA) and Efficient Global Self-Attention(EGSA). The former is designed to capture local relational information among pixels within spatially adjacent regions, while the latter focuses on modeling long-range dependencies and similarities among changed regions across the entire image. The detailed architecture of LGFB is illustrated in Figure 3. Structurally, LGFB is a Transformer-like module. However, conventional self-attention mechanisms typically incur substantial computational and parametric overhead. To address this, we adopt a more efficient design, which is described in detail below.
Token Mixer & Channel Mixer. Inspired by the RepViT [14], we decompose the attention computation into two stages: a token mixer and a channel mixer, where the former is responsible for aggregating spatial information and the latter focuses on inter-channel interaction. Specifically, in both the SWSA and EGSA modules, we adopt depthwise separable convolutions(DWConv) for linear projections. While this projection lacks explicit interaction across channels, our subsequent experiments demonstrate that this design substantially reduces the number of parameters without compromising model performance. The resulting projected matrices are then fed into the attention mechanism, enabling spatial information exchange with minimal parameter and computational overhead—serving as the token mixer. This process can be formally expressed as:
$Q=DWConv_{1}(I),$
$K=DWConv_{2}(I),$
$V=DWConv_{3}(I),$
@@
Here, $I$ represents the input to the SWSA and EGSA modules. Subsequently, a multilayer perceptron (MLP) composed of two $1\times 1$ convolutional layers is employed to perform information exchange along the channel dimension, serving as the channel mixer. This process can be formally expressed as:
@@
Sliding Window Self-Attention.To enable the model to focus on local pixel-level relationships, we constrain the attention computation within a sliding window. A stride parameter $s$ is introduced to control how the window moves across the entire feature map, producing the final attention output. Both the window size $w$ and the stride $s$ are treated as hyperparameters. When the window size is equal to the stride ($w=s$), the feature map is partitioned into non-overlapping patches, and attention is computed independently within each patch. In contrast, when the window size is larger than the stride ($w>s$), only a portion of the computed attention within each window is retained as a patch, resulting in overlapping receptive fields where the patch size is $s\times s$. In scenarios where change regions exhibit complex shapes, using a larger window with overlapping receptive fields enhances spatial continuity across neighboring regions. Conversely, for regions with low shape complexity, such as regular building changes, setting the window size equal to the stride is sufficient to capture local patterns. It avoids unnecessary contextual redundancy while reducing additional computational overhead.
@@
Efficient Global Self-Attention.To capture global relationships among change regions and further mitigate the spatial discontinuity introduced by patch-wise processing in SWSA, we introduce Efficient Global Self-Attention (EGSA) following SWSA. In EGSA, we first apply linear projections to the input $D_{L}$ to obtain the query $Q_{G}$, key $K_{G}$, and value $V_{G}$ matrices. The $Q_{G}$ matrix retains the same spatial resolution as the input to preserve detailed spatial information, while the $K_{G}$ and $V_{G}$ matrices are downsampled to a lower resolution. This downsampling serves to compress global information, effectively reducing both the parameter count and computational cost. Specifically, the $K_{G}$ and $V_{G}$ matrices are obtained via convolution operations, where both the kernel size and stride are set to match the patch size used in SWSA. This design ensures that each token in $K_{G}$ and $V_{G}$ corresponds to a patch from the input $D_{L}$. We then compute the attention weights between each pixel in $Q_{G}$ and all patch-level tokens in $K_{G}$, enabling the model to aggregate long-range contextual information. The final output feature map is generated by applying the computed attention weights to the $V_{G}$ matrix.
@@
Hierarchical Feature Integration.The lightweight decoder takes as input the multi-scale difference features $D_{i}(i=1,2,3)$ generated from three hierarchical stages. Among them, higher-level difference maps exhibit lower spatial resolution but encode richer semantic information about change regions, while lower-level difference maps preserve finer spatial details such as object boundaries but tend to be less reliable in distinguishing actual changes. To leverage the complementary advantages of different levels, we adopt a progressive fusion strategy. As illustrated on the left side of Figure 2, each higher-level difference map is processed by the LGFB, and then fused with the corresponding lower-level difference map via element-wise addition. A depthwise separable convolution is then applied to refine the fused features, achieving a local adjustment. However, this local fusion alone is not sufficient to capture long-range dependencies across spatially distant regions. Therefore, we further incorporate EGSA after the local fusion to perform global refinement and enhance the fusion quality.
@@
Figure 4: Changed region characteristics in four datasets. For each dataset, samples are divided by the number of change regions ($\leq$4 on the left, $>$4 on the right). The x-axis denotes shape complexity (perimeter/area); the y-axis shows area ratio for the left plots and region similarity (variance of shape complexity) for the right plots.
@@ 4.1 Datasets and Experiment Settings
Dataset Preparation and Analysis. To validate the effectiveness of FlickCD, we conduct experiments on four widely-used change detection datasets: SYSU [27], WHU [28], LEVIR+ [29], and CDD [30]. To better interpret the experimental results, we perform a statistical analysis on the characteristics of change regions within each dataset, as shown in Figure 4. We categorize the samples in each dataset into two groups based on the number of change regions: few-region samples and many-region samples, using a threshold of 4 change regions. For few-region samples, we analyze the area proportion of change regions and their shape complexity. For many-region samples, we focus on the inter-region dissimilarity and shape complexity. The analysis reveals that:
  • SYSU primarily consists of samples with a small number of change regions, where most changes occupy a large area and exhibit relatively high shape complexity.
  • CDD predominantly contains samples with many change regions, characterized by high variability in inter-region similarity and complex region shapes.
  • WHU and LEVIR+ focus mainly on building-related changes, leading to a large number of multi-region samples. In these datasets, the change regions tend to exhibit high inter-region similarity and low shape complexity, reflecting the relatively regular structure of architectural changes.
Experiment Settings. To accommodate the varying characteristics of change regions across datasets, we configure the hyperparameters of the Sliding-Window Self-Attention (SWSA) module, namely the window size $w$ and stride $s$ individually for each dataset. Given that higher-level feature maps have lower spatial resolution, we assign smaller window sizes at deeper layers accordingly.For the SYSU and CDD datasets, the change regions tend to exhibit higher shape complexity, requiring a larger receptive field to improve the accuracy of feature extraction. Accordingly, we set the window sizes to 8,8,16 with corresponding strides of 4,4,8 for these datasets. In contrast, the change regions in WHU and LEVIR+ are generally more regular and structurally simple. For these datasets, we adopt an equal window size and stride configuration to reduce computational overhead without sacrificing performance. Specifically, we set the parameters to 4,4,8 for WHU and 4,8,8 for LEVIR+. For model training, the batch size is set to 32. The learning rate is set to 0.0005 for all datasets except WHU, which uses a lower learning rate of 0.0002. We train the model for 100 epochs on the SYSU and WHU datasets, 200 epochs on LEVIR+, and 250 epochs on CDD.
@@
4.2 Overall Comparison
Comparison Baselines. To demonstrate the trade-off advantage of FlickCD in change detection tasks, we compare it against both lightweight and high-performance CD models. The lightweight models include the FC series [15] (FC-EF, FC-Siam-Diff, FC-Siam-Conc), TinyCD [10], and USSFC-Net [9], all of which are based on convolutional neural networks (CNNs). The high-performance models include RCTNet [7](CNN-based), ChangeFormer [12] and BiFA [31](Transformer-based), and ChangeMamba [8](based on the recent Mamba architecture). Through this comparison, we aim to highlight the efficiency advantage of FlickCD over high-performance models, and its performance advantage over lightweight models.
Quantitative Results. We primarily compare the F1 scores of different models across four datasets, as this metric provides a comprehensive evaluation of model performance. As shown in Table 1 and 2, it is evident that models with fewer parameters than FlickCD perform significantly worse across all datasets. Compared with the models that have more parameters, FlickCD achieves the highest performance in SYSU and WHU datasets, and maintains performance on par with the best-performing model ChangeMamba in CDD and LEVIR+ datasets. Specifically, FlickCD achieves an F1-score of 83.97% on the SYSU dataset and 94.81% on the WHU dataset, surpassing the second-best model by 0.14% and 0.25%, respectively. Although FlickCD does not outperform ChangeMamba in terms of F1 score on the CDD and LEVIR+ datasets, it maintains a comparable performance level and ranks second among the selected models.
Qualitative Results. To gain a more intuitive and detailed understanding of model performance in CD tasks, we select samples from the SYSU, WHU, and LEVIR+ datasets and compare the output of five resource-intensive models with FlickCD. As shown in Figures 5 and 6, FlickCD demonstrates superior recognition performance on challenging change regions. Specifically, for Figure 6 and sample (3) in Figures 5, some color variations caused by illumination differences are actually irrelevant changes. However, the comparison models mistakenly classify them as significant change areas, while FlickCD correctly identifies and filters out these irrelevant differences. Furthermore, although FlickCD achieves a slightly lower F1-score than ChangeMamba on the LEVIR+ dataset, it shows better extraction performance in cases involving multiple similar change regions, such as sample (4) in Figure 5. This visual comparison highlights FlickCD’s stronger ability to distinguish irrelevant differences and more accurate boundary extraction for change regions.
Model Efficiency Analysis To compare the computational efficiency and resource requirements of different models, we report the number of parameters, FLOPs, throughput, and inference time for each model. Specifically, the inference time refers to the time required to process a single sample, while the throughput is calculated by dividing the number of samples in one epoch by the total runtime of that epoch. The corresponding results are presented in Table 1 and 3. The FC series, TinyCD, and USSFC-Net significantly reduce the parameter count at the cost of considerable performance degradation. However, FC models and USSFC-Net still incur relatively high computational costs. Among the high-performance models, although they achieve competitive results across various datasets, their resource demands are substantial. For example, ChangeMamba, which demonstrates a good balance of accuracy across datasets, requires 53.99M parameters and 31.3 ms inference time per sample. In contrast, FlickCD achieves comparable performance with only 1.89M parameters,approximately 3.5% of ChangeMamba, and a lower inference time of 9.2 ms, which is less than one-third of that of ChangeMamba. In addition, FlickCD achieves a throughput that is 2 to 3 times higher than that of the high-performance models. These results clearly demonstrate the efficiency of FlickCD in terms of parameter count, computation cost, and latency, while maintaining strong performance.
@@ 4.3 Ablation Study
Effectiveness of EDM To validate the effectiveness of the Enhanced Difference Module(EDM) in FlickCD, we first conducted a visual ablation study, with the results shown in Figure 6. On the SYSU dataset, FlickCD exhibits a stronger ability to distinguish between irrelevant and meaningful changes. To confirm that this capability stems from the EDM, we remove EDM and retrained the model, using a direct Manhattan distance to compute the difference features between the two temporal inputs. As illustrated in the results, the model fails to distinguish irrelevant changes from significant ones without EDM, similar to other baseline methods. In addition to the visual comparison, we also performed quantitative ablation experiments across all datasets, as shown in Table 4. By integrating the EDM into the base architecture, we observe consistent improvements in F1-scores across all datasets, with the most significant gain of 0.65% on SYSU. These findings from both visual and quantitative perspectives demonstrate the effectiveness of the EDM module in enhancing the model’s discriminative capability.
Figure 6: Visualization results on the SYSU dataset, showcasing the outputs of other models, FlickCD with EDM, and FlickCD without EDM. The comparison highlights FlickCD’s superior difference discrimination capability and validates the effectiveness of the EDM.
@@
Effectiveness of SWSA and EGSA. To further verify the effectiveness of Sliding-Window Self-Attention(SWSA) and Efficient Global Self-Attention(EGSA) in analyzing difference information, we conducted ablation experiments involving these two modules on all datasets. The results are summarized in Table 4, revealing two distinct trends:
@@
These findings, when correlated with the dataset characteristics illustrated in Figure 4, provide further insights. Both WHU and LEVIR+ datasets contain a large number of multi-instance change regions, with high similarity among regions and low shape complexity. In such cases, omitting one attention scale (local or global) leads to a greater loss than the gain achieved by using only one. This suggests that both local and global contextual analyses are essential for effectively capturing meaningful change information in datasets characterized by numerous, similar, and relatively simple change regions.
@@
  • Employing depthwise separable convolutions for linear projections in both EGSA and SWSA.
@@
To assess the potential trade-offs introduced by these strategies, we conducted ablation experiments, with the results summarized in Table 5. The findings indicate that both strategies reduce the model’s parameter count by nearly half, yet no performance degradation was observed across the four datasets. This suggests that the information loss induced by these strategies is non-critical. Specifically, while depthwise separable convolutions discard inter-channel interactions, this limitation is compensated by the subsequent MLP module, which facilitates cross-channel communication. Furthermore, the third encoder stage already provides sufficient semantic abstraction for change region extraction, and incorporating the fourth stage does not further improve performance.

### v2

@@ 3.3 Local-Global Fusion Block
After obtaining the difference map $D_{i}$ between the two feature maps, it is essential to further capture the relational information among the changed regions. To this end, we extract relationship features at both local and global scales, motivated by the following observations:
  • On the local scale, when a specific region is surrounded by other changed areas, it is highly likely that this region has also undergone change. Modeling such local relationships helps mitigate the presence of holes or fragmentation within detected changed regions.
  • On the global scale, by leveraging global contextual information, the model can enhance overall feature representations, thereby increasing the discriminability between changed and unchanged regions and facilitating more accurate delineation of change areas.
To achieve the aforementioned objectives, we propose the Local-Global Fusion Block (LGFB). This module consists of two key components: Sliding Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA). The former is designed to capture local relational information among pixels within spatially adjacent regions, while the latter focuses on modeling long-range dependencies and enhancing feature representation. The detailed architecture of LGFB is illustrated in Fig. 3. Structurally, LGFB is a Transformer-like module. However, traditional self-attention mechanisms are often computationally and parameter-intensive. To address this, we adopt a more efficient design, which is described in detail below.
Token Mixer & Channel Mixer. Inspired by the RepViT [14], we decompose the attention computation into two stages: a token mixer and a channel mixer, where the former is responsible for aggregating spatial information and the latter focuses on inter-channel interaction. Specifically, in both the SWSA and EGSA modules, we adopt Depth-wise Separable Convolution (DWConv) for linear projections. The projected matrices are then fed into the attention mechanism to efficiently exchange spatial information, acting as a lightweight token mixer. This process can be formally expressed as:
$Q,K,V=DWConv_{1}(I),DWConv_{2}(I),DWConv_{3}(I),$
@@
Here, $I$ represents the input to the SWSA and EGSA modules. Subsequently, a multi-layer perceptron (MLP) composed of two $1\times 1$ convolutional layers is employed to perform information exchange along the channel dimension, serving as the channel mixer. This process can be formally expressed as:
@@
Sliding Window Self-Attention.To enable the model to focus on local pixel-level relationships, we constrain the attention computation within a sliding window. A stride parameter $s$ is introduced to control how the window moves across the entire feature map, producing the final attention output. Both the window size $w$ and the stride $s$ are treated as hyperparameters. When the window size is equal to the stride ($w=s$), the feature map is partitioned into non-overlapping patches, and attention is computed independently within each patch. In contrast, when the window size is larger than the stride ($w>s$), only a portion of the computed attention within each window is retained as a patch, resulting in overlapping receptive fields where the patch size is $s\times s$. In scenarios where changed regions exhibit complex shapes, using a larger window with overlapping receptive fields enhances spatial continuity across neighboring regions. Conversely, for regions with low shape complexity, such as regular building changes, setting the window size equal to the stride is sufficient to capture local patterns. It avoids unnecessary contextual redundancy while reducing additional computational overhead.
@@
Efficient Global Self-Attention.To capture global relationships among changed regions and further mitigate the spatial discontinuity introduced by patch-wise processing in SWSA, we introduce Efficient Global Self-Attention (EGSA) following SWSA. In EGSA, we first apply linear projections to the input $D_{L}$ to obtain the query $Q_{G}$, key $K_{G}$, and value $V_{G}$ matrices. The $Q_{G}$ matrix retains the same spatial resolution as the input to preserve detailed spatial information, while the $K_{G}$ and $V_{G}$ matrices are downsampled to a lower resolution. This downsampling serves to compress global information, effectively reducing computational cost. Specifically, the $K_{G}$ and $V_{G}$ matrices are obtained via convolution operations, where both the kernel size and stride are set to match the patch size used in SWSA. This design ensures that each token in $K_{G}$ and $V_{G}$ corresponds to a patch from the input $D_{L}$. We then compute the attention weights between each pixel in $Q_{G}$ and all patch-level tokens in $K_{G}$, enabling the model to aggregate long-range contextual information. The final output feature map is generated by applying the computed attention weights to the $V_{G}$ matrix.
@@
Hierarchical Feature Integration.The lightweight decoder takes the multi-scale difference features $D_{i},i\in\{1,2,3\}$ generated from three hierarchical stages, as input. Among them, higher-level difference maps have lower spatial resolution but capture richer semantic information about changed regions, whereas lower-level maps retain finer spatial details, such as object boundaries, but are less reliable in distinguishing actual changes. To leverage the complementary advantages of different levels, we adopt a progressive fusion strategy. As illustrated in Fig. 2(a), each higher-level difference map is processed by the LGFB, and then fused with the corresponding lower-level difference map via element-wise addition. A Depth-wise Separable Convolution is then applied to refine the fused features, achieving a local adjustment. However, this local fusion alone is not sufficient to capture long-range dependencies across spatially distant regions. Therefore, we further incorporate EGSA after the local fusion to perform global refinement and enhance the fusion quality.
@@
Figure 4: Changed region characteristics in four datasets. For each dataset, samples are divided by the number of changed regions ($\leq$4 on the left, >>4 on the right). The x-axis denotes shape complexity (perimeter/area); the y-axis shows area ratio for the left plots and region similarity (variance of shape complexity) for the right plots.
@@ 4.1 Datasets and Experiment Settings
To validate the effectiveness of FlickCD, we conduct experiments on four widely-used change detection datasets: SYSU [27], WHU [28], LEVIR+ [29], and CDD [30]. To better interpret the experimental results, we perform a statistical analysis on the characteristics of changed regions within each dataset, as shown in Fig. 4. We categorize the samples in each dataset into two groups based on the number of changed regions: few-region samples and many-region samples, using a threshold of 4 changed regions. For few-region samples, we analyze the area proportion of changed regions and their shape complexity. For many-region samples, we focus on the inter-region dissimilarity and shape complexity. The analysis reveals that:
  • SYSU primarily consists of samples with a small number of changed regions, where most changes occupy a large area and exhibit relatively high shape complexity.
  • CDD mainly consists of samples with numerous changed regions, which show large variation in similarity between regions and exhibit complex shapes.
  • WHU and LEVIR+ focus mainly on building-related changes, leading to a large number of multi-region samples. In these datasets, the changed regions tend to exhibit high inter-region similarity and low shape complexity, reflecting the relatively regular structure of architectural changes.
Experiment Settings. To accommodate the varying characteristics of changed regions across datasets, we configure the hyperparameters of the Sliding-Window Self-Attention (SWSA) module, namely the window size $w$ and stride $s$ individually for each dataset. Given that higher-level feature maps have lower spatial resolution, we assign smaller window sizes at deeper layers accordingly.For the SYSU and CDD datasets, the changed regions tend to exhibit higher shape complexity, requiring a larger receptive field to improve the accuracy of feature extraction. Accordingly, we set the window sizes to 8,8,16 with corresponding strides of 4,4,8 for these datasets. In contrast, the changed regions in WHU and LEVIR+ are generally more regular and structurally simple. For these datasets, we adopt smaller window sizes and strides configuration to reduce computation. Specifically, we set the parameters to 4,4,8 for WHU and 4,8,8 for LEVIR+. For model training, the batch size is set to 32. The learning rate is set to 0.0005 for all datasets except WHU, which uses a lower learning rate of 0.0002. We train the model for 100 epochs on the SYSU and WHU datasets, 200 epochs on LEVIR+, and 250 epochs on CDD.
@@
4.2 Overall Comparison
To demonstrate the trade-off advantage of FlickCD in CD tasks, we compare it against both lightweight and high-performance CD models. The lightweight models include the FC series (FC-EF, FC-Siam-Diff, FC-Siam-Conc) [15], TinyCD [10], and USSFC-Net [9]. The high-performance models include RCTNet [7], ChangeFormer [12], BiFA [31], and ChangeMamba [8]. Through this comparison, we aim to highlight the efficiency advantage of FlickCD over high-performance models, and its performance advantage over lightweight models.
Quantitative Results. We primarily compare the F1 scores of different models across four datasets, as this metric provides a comprehensive evaluation of model performance. As shown in Table 1 and 2, it is evident that models with fewer parameters than FlickCD perform significantly worse across all datasets. Compared with the models that have more parameters, FlickCD achieves the highest performance in SYSU and WHU datasets, and maintains performance on par with the best-performing model ChangeMamba in CDD and LEVIR+ datasets. Although FlickCD does not outperform ChangeMamba in terms of F1 score on the CDD and LEVIR+ datasets, it maintains a comparable performance level and ranks second among the selected models.
Qualitative Results. To gain a more intuitive and detailed understanding of model performance in CD tasks, we select samples from the SYSU, WHU, and LEVIR+ datasets and compare the output of five resource-intensive models with FlickCD. As shown in Fig. 5 and 6, FlickCD demonstrates superior recognition performance on challenging changed regions. Specifically, for Fig. 5(3) and Fig. 6, some color variations caused by illumination differences are actually irrelevant changes. However, the comparison models mistakenly classify them as significant changes, while FlickCD identifies and filters out these irrelevant differences. Furthermore, although FlickCD achieves a slightly lower F1-score than ChangeMamba on the LEVIR+ dataset, it shows better extraction performance in cases involving multiple similar changed regions, such as Fig. 5(4). This visual comparison highlights FlickCD’s stronger ability to distinguish irrelevant differences and more accurate extraction for changed regions.
@@ 4.2 Overall Comparison
Model Efficiency Analysis To compare the computational efficiency and resource requirements of different models, we report the number of parameters, FLOPs, throughput, and inference time for each model. Specifically, the inference time refers to the time required to process a single sample, while the throughput is calculated by dividing the number of samples in one epoch by the total runtime of that epoch. The corresponding results are presented in Table 1 and 3. The FC series, TinyCD, and USSFC-Net significantly reduce the parameter count at the cost of considerable performance degradation. However, the FC series and USSFC-Net still have relatively high FLOPs values. Among the high-performance models, although they achieve competitive results across various datasets, their resource demands are substantial. For example, ChangeMamba, which demonstrates a good balance of accuracy across datasets, requires 53.99M parameters and 31.3 ms inference time per sample. In contrast, FlickCD achieves comparable performance with only 1.89M parameters, approximately 3.5% of ChangeMamba, and a lower inference time of 9.2 ms, which is less than one-third of that of ChangeMamba. In addition, FlickCD achieves a throughput that is 2 to 3 times higher than that of the high-performance models. These results clearly demonstrate the efficiency of FlickCD in terms of parameter count, computation cost, and latency, while maintaining strong performance.
@@ 4.3 Ablation Study
Effectiveness of EDM To validate the effectiveness of the Enhanced Difference Module (EDM) in FlickCD, we first conducted a visual ablation study, with the results shown in Fig. 6. On the SYSU dataset, FlickCD exhibits a stronger ability to distinguish between irrelevant and critical changes. To confirm that this capability stems from the EDM, we removed EDM and retrained the model, obtaining the difference features between the two inputs by subtracting their feature maps and taking the absolute value. As illustrated in the results, the model fails to distinguish irrelevant changes from significant ones without EDM, similar to other baseline methods. In addition to the visual comparison, we also performed quantitative ablation experiments across all datasets, as shown in Table 4. By integrating the EDM into the base architecture, we observe consistent improvements in F1-scores across all datasets, with the most significant gain of 0.65% on SYSU. These findings from both visual and quantitative perspectives demonstrate the effectiveness of the EDM in enhancing the model’s discriminative capability.
Figure 6: Visualization results on the SYSU dataset, showcasing the outputs of other models, FlickCD with EDM, and FlickCD without EDM.
@@
Effectiveness of SWSA and EGSA. To verify the effectiveness of Sliding-Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA) in analyzing difference information, we conducted ablation experiments involving these two modules on all datasets. The results are summarized in Table 4, revealing two distinct trends:
@@
These findings, when correlated with the dataset characteristics illustrated in Fig. 4, provide further insights. Both WHU and LEVIR+ datasets contain a large number of multi-instance changed regions, with high similarity among regions and low shape complexity. In such cases, omitting one attention scale (local or global) leads to a greater loss than the gain achieved by using only one. This suggests that both local and global contextual analyses are essential for effectively capturing meaningful changed information in datasets characterized by numerous, similar, and relatively simple changed regions.
Figure 7: Visualization of intermediate feature maps and attention heatmaps: (a) Difference map before EDM; (b) Attention heatmap of EDM; (c) Feature map after EDM; (d) Feature map after SWSA; (e) Feature map after EGSA; (f) Final result.
Visualization of Intermediate Features. To further demonstrate the effectiveness of each module, we visualize several intermediate feature maps and attention heatmaps, as shown in Fig. 7. The difference map before EDM processing fails to accurately detect the changed regions. In contrast, the attention heatmap generated by EDM reallocates the weights of different regions, resulting in a more precise segmentation of the changed areas (Fig. 7(c)). Subsequently, SWSA refines the delineation of local changed regions, while EGSA further enhances the global difference information. Finally, the model outputs an accurate binary change detection map.
@@
  • Employing depth-wise separable convolutions for linear projections in both EGSA and SWSA.
@@
To assess the potential trade-offs introduced by these strategies, we conducted ablation experiments, with the results summarized in Table 5. The findings indicate that both strategies reduce the model’s parameter count by nearly half, yet no performance degradation was observed across the four datasets. This suggests that the information loss induced by these strategies is non-critical. Specifically, while depth-wise separable convolutions discard inter-channel interactions, this limitation is compensated by the subsequent MLP, which facilitates cross-channel communication. In addition, the third encoder stage already provides sufficient semantic information for changed region extraction, and incorporating the fourth stage does not further improve performance.